OpenStudy (anonymous):

what is an equation of a line parallel to y=7x+5 with an x-intercept of 6

OpenStudy (amriju):

ur given eqn is in the form y=mx+c...so m(slope) is 7...now for a line parallel to ur eqn..the slope remains same...so u havethe eqn of the line as y=7x+c...where c is the a constant to be determine...Now X intercept( thats where the line cuts the X axis) is 6..so at that point where the line cuts the X axis..the y co-ordinate should be 0... so the line passes through..(6,0)..so put (6,0) in place of x any Y...the eqn should be satisfied...so..0=42+c..so c=-42...hence eqn of line y=7x-42

OpenStudy (luigi0210):

^Question answered, what more do you need? btw Welcome to openstudy :)

OpenStudy (nurali):

First, get the equation into y =mx + b form. y=7x+5 Here, we know the slope (m) is 7, so the slope of our new line is also 7. We also know a point on the new line (6, 0). So, let's figure out the y intercept. y = 7x + b 0 = 7(6) + b 0 = 42 + b b = -42 So, the equation is y = 7x - 42. In ax + by = c form, just rearrange. y = 7x - 42 7x-y=42

OpenStudy (nurali):

Welcome to Openstudy @mizzougirl76